Increaing a number by five and then doubling the outcome is the same same as decresing the number by twelve
Answer by rapaljer(4551)

2(x+5)= x-12
2x + 10 = x-12

Subtract x from each side:
2x -x + 10 = x-x - 12
x+10 = -12

Subtract 10 from each side:
x+10-10= -12-10
x= -22

Check: Follow the instructions of the problem.
Increase the number by 5, which would be -17, then double this would give you -34.

Decrease the same number -22 by 12 also gives you -34.
It checks!!

R^2 at SCC